Homework 3, February 5

Math 402 – Spring Semester, 1990 Second Course in Abstract Algebra, II Prof. Daniel Grayson

2:00 PM MWF, Section F1, 243 Altgeld Hall

These problems are due Wednesday, February 14. Do p. 365, problems 1, (2), 3, (5), (6), 7, 8, (9), 10, 13; and p. 369, problems 4, (5),

(6), (7), 9, 10.

The exercises in parentheses are to be turned in only if you are uncertain whether you got the correct answer.